OpenStudy (oreolover4876):

R, A, and T are midsegments. The perimeter of triangle RAT is 15cm. what is the perimeter of triangle LOB?

OpenStudy (anonymous):

Please draw a figure, we cannot see LOB and RAT :)

OpenStudy (oreolover4876):



OpenStudy (anonymous):

Perimeter of RAT = RA + AT + RT Because A is mid-segment of LB and RA // BO, we have: \[\frac{ RA }{ BO } = \frac{LA}{LB} = \frac{1}{2}\] or BO = 2*RA. Similarly, LO = 2*AT ; LB = 2*RT Perimeter of LOB = LO + BO + LB = 2* (RA + AT + RT) = 2*15 = 30(cm)